hiya:

Given: cos x= square root of 2/2 What is sin x and tan x? Explain your steps in complete sentences. Can someone please help me fast, i'm taking a test rn and I only have like 15 minutes left. Please HELP MEE!!!

hiya:

@justus

justus:

First we need to know what hypotenuse, Adjacent, and Opposite sides are.

justus:

sin x = opposite / hypotenuse.

justus:

adj = square root (hypotenuse^2 - opposite^2) adj = square root (2^2 - square root(2)^2) adj= square root (4 - 2) adj= square root (2)

hiya:

this is what i currently did O= H=2 Adj=?

1 attachment
justus:

Hmm this is the pythagorean theorem.

hiya:

yes is it not right?

justus:

This is not what you asked.

hiya:

oh well how do you solve it?

justus:

the question you asked?

justus:

I am almost finished.

hiya:

ok please try to hurry up i'm in a life or death situation rn

justus:

To get your answer cos x = adjacent / hypotenuse tan x = opposite / adjacent add the adjacent and hypotenuse and the opposite side and adjacent

justus:

I mean plug them in

hiya:

i don't understand anything rn can you please do it for me? i would really appreciate if you could

justus:

I would love to but I am not allowed to give direct answers.

hiya:

I literally have like 7 minutes left

justus:

can you try and I will tell you if its correct?

hiya:

ok one sec

justus:

sure

hiya:

is this right? cos x = 2 / 4=1/2

hiya:

i don't know how to do tanx

justus:

very close, it is actually: cos x = square root (2) / 2 and tan x = 1

justus:

I explained above how to get the answer

hiya:

oh okay but can you help me put them in numbers please

hiya:

omg i was supposed to find the sin x not the cos x

justus:

huh?

justus:

oh

hiya:

is the sin x and cos x the same?

justus:

Sin(x) = opposite over hypotenuse

justus:

Cos(x) = adjacent over hypotenuse

hiya:

how do you find the sinx? and tan x?

justus:

Opposite is the square root of 2 and the hypotenuse = 2 H^2 = O^2 + A^2 2^2 = (square root of 2)^2 + A^2 A^2 = 4 - 2 A = square root of 2 Cos(x) = square root of 2 over 2 Tan(x) = square root of 2 over square root of 2 = 1

justus:

are there options btw?

hiya:

no it's an open response and the cos x was already given. and I literally did the same thing

justus:

Oh

hiya:

please help me fast i also have an another question to do and i don't have enought time

justus:

oh okay just a second

hiya:

ok

hiya:

can you also help me do this too? Using complete sentences, explain the key features of the graph of the sine function.

hiya:

this is what i did before for the first question

1 attachment
hiya:

@justus are you still there?

justus:

yes im trying to see how to solve it

hiya:

oh ok please try to hurry up

hiya:

@justus u still there?

hiya:

@justus just tell me how to find tan x

justus:

tan x its sin divided by its cosine: tan x = sin x cos x .

hiya:

I answere d the first question Can you please help me answer this? Using complete sentences, explain the key features of the graph of the sine function.

justus:

hopefully this helps you figure it out https://www.mathsisfun.com/sine-cosine-tangent.html